Bob is 24 years Younger than his father. In 2 years he will be half his father's age. How old is each now?
what i understand is that B= Bob's age, and B+24= Fathers age and i don't understand how we are supposed to write out the 2 years part of the equation.

Answer by scott8148(6628)

F = B + 24

(F + 2) / 2 = B + 2